A cylinder has a radius of 8 inches and a volume of 2,009.6 cubic inches.
What is the height of the cylinder?

Answer by mananth(16946)

Volume of cylinder = pi*r^2*h
Volume = 2009.6 cu.in
radius= 8 inches

Volume = pi* 8 ^2 * h 3.14
2009.6 = 200.96 h

h= 2009.6 / 200.96
h= 10 inches